### v1

# Multilingual large language models leak human stereotypes across language boundaries

Abstract.
Multilingual large language models have been increasingly popular for their proficiency in comprehending and generating text across various languages. Previous research has shown that the presence of stereotypes and biases in monolingual large language models can be attributed to the nature of their training data, which is collected from humans and reflects societal biases. Multilingual language models undergo the same training procedure as monolingual ones, albeit with training data sourced from various languages. This raises the question: do stereotypes present in one social context leak across languages within the model? In our work, we first define the term “stereotype leakage” and propose a framework for its measurement. With this framework, we investigate how stereotypical associations leak across four languages: English, Russian, Chinese, and Hindi. To quantify the stereotype leakage, we employ an approach from social psychology, measuring stereotypes via group-trait associations. We evaluate human stereotypes and stereotypical associations manifested in multilingual large language models such as mBERT, mT5, and ChatGPT. Our findings show a noticeable leakage of positive, negative, and non-polar associations across all languages. Notably, Hindi within multilingual models appears to be the most susceptible to influence from other languages, while Chinese is the least. Additionally, ChatGPT exhibits a better alignment with human scores than other models.
WARNING: This paper contains model outputs which could be offensive in nature.
1. Introduction
Cultural stereotypes about social groups can be transmitted based on how these social groups are represented, treated, and discussed within each culture (Martinez et al., 2021; Lamer et al., 2022; Rhodes et al., 2012). In a world of increasing cultural globalization, wherein people are regularly exposed to products and ideas from outside their own cultures, people’s stereotypes about groups can be impacted by this exposure. For instance, blackface is characterized as one of America’s first cultural exports, as the performance of American minstrelsy shows in different countries popularized racist depictions of Black Americans within those other cultures (THELWELL, 2020).
Recently, the deployment of large language models has the potential to exacerbate the issue. Large language models are becoming increasingly language-agnostic. For instance, models like ChatGPT(OpenAI111https://openai.com/chatgpt, we use GPT text-davinci-003 version(Ouyang et al., 2022)) and mBART (Lin et al., 2022) can operate without being restricted to a specific language, handling input and output in multiple languages simultaneously. This thus gives rising opportunities for what we refer to as stereotype leakage, or the transmission of stereotypes from one culture to another.
Stereotype leakage within large language models may export harmful stereotypes across cultures and reinforce Anglocentricism222Anglocentrism is the practice of viewing and interpreting the world from an English-speaking perspective with the prioritization of English culture, language, and values. Anglocentrism can lead to biases and neglect of global perspectives and experiences.. Previous works (e.g., Goldstein et al., 2023; Weidinger et al., 2021) have highlighted the potential for language model outputs to change users’ perceptions and behaviors. Stereotype leakage from large language models may further entrench existing stereotypes among model users, as well as create new stereotypes that have transferred from a different language. Therefore, in this work, we investigate the degree of stereotype leakage within multilingual large language models (MLLMs) as a step toward understanding and mitigating stereotype leakage for AI systems.
Large language models are currently the backbone of many natural language processing (NLP) models. MLLMs are language models pre-trained with a large amount of data from multiple languages so that they can process NLP tasks in various languages as well as cross-lingual tasks. Recent MLLMs, such as GPT models (Brown et al., 2020; OpenAI, 2023) designed for standalone applications and models such as mBERT (Müller et al., 2020), XLM (Lample and Conneau, 2019), mT5 (Xue et al., 2020), mBART (Lin et al., 2022), intended for use as back-end tools, show satisfactory performance on NLP tasks across around 100 languages. One major advantage of such models is that low-resource languages (languages with less training data) can benefit from high-resource languages through shared vocabulary (Lample and Conneau, 2019) and structural similarities (word-ordering or word-frequency) (K et al., 2020).
Large language models are trained on existing language data, and even monolingual language models have been demonstrated to replicate stereotypical associations present in the training data. (Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022). Thus, with the shared knowledge between languages in MLLMs, it is likely that stereotypes may also leak between languages. Stereotypes are abstract and over-generalized pictures drawn about people based on their group membership, and these perceptions can be specific to each culture. Though LLMs are trained on language-based data rather than culture-based data, languages reflect the stereotypes associated with the cultures they represent. Thus, for the purpose of studying stereotypes in MLLMs, we divide the world according to languages, with the understanding that a single language may reflect multiple cultures. Previously, many works have examined Western stereotypes in English language models (e.g. Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022), whereas limited works have attempted to assess stereotypes in multilingual language models (e.g. Kaneko et al., 2022; Levy et al., 2023; Câmara et al., 2022) due to the complexity of stereotypes manifested in various cultures, limited resources, and Anglocentric norms (Talat et al., 2022).
In this paper, we aim to investigate the existence of stereotype leakage in MLLMs, which we define as the effect of stereotypical word associations in MLLMs of one language impacted by stereotypes from other languages. We conduct a human study to collect human stereotypes, adopt word association measurement approaches from previous works (Cao et al., 2022; Kurita et al., 2019) to measure stereotypical associations in MLLMs and analyze the strength and nature of stereotype leakage across different languages both quantitatively and qualitatively.
To test our hypothesis that there are significant stereotypes leaked across languages in MLLMs, we sample four languages: English, Russian, Chinese, and Hindi. We pick languages that come from the Indo-European and Sino-Tibetan language families, ranging from high (English) to low-resource (Hindi) languages333 High-resource languages are languages that have more training data available, while low-resource languages have less.. We measure the degree of stereotype leakage between the four languages in three MLLMs — mBERT, mT5, and ChatGPT444Both the code and the dataset, along with a datasheet (Gebru et al., 2018), are available under a MIT licence at: https://github.com/AnnaSou/Stereotype_Leakage.. Both mBERT and mT5 are back-end MLLMs. MT5 has better multilingual performance than mBERT, whereas mBERT has more comparable monolingual BERT models for the four languages. ChatGPT is one of the state-of-the-art MLLMs that has been popularly deployed to users. With these, we examine the impact of human stereotypes from different languages on stereotypical associations in MLLMs.
2. Measuring Stereotype Leakage in MLLMs
For each language, we aim to assess the degree of stereotype leakage from the other languages to this target language in MLLMs. Specifically, we measure the effect of human stereotypes from all four languages () on the target language’s MLLM stereotypical association (), as shown in Equation 1. For mBERT, we also measure the impact of the stereotypical association from the target monolingual model (). We use a mixed-effect model to fit the formula and calculate the effect. If the coefficient of a variable is positive and has a p-value of less than , then the variable has a significant effect on . If there are significant effects from the non-target language’s human stereotypes, then there are potential stereotype leakages from this language to the target language.
In the following section, we discuss how we measured each of the variables.
2.1. Stereotype Measurement
In this paper, we measure stereotypes through group-trait associations with traits from the Agency Beliefs Communion (ABC) model of stereotype content (Koch et al., 2020). The model consists of 16 trait pairs (each pair represents two polarities) that are designed to characterize group stereotypes along the dimensions of agency/socioeconomic success, conservative–progressive beliefs, and communion, as listed in Table 1.
If a group (e.g. “immigrant”, “Asian person”) has a high degree of association with a trait (e.g. religious, confident), then we consider that trait a stereotype of the group. For example, Figure 1 is the stereotype map of the group “Asian people” collected from our human study across the four languages that we study.
For the groups, we picked groups listed in Table 2: shared groups with shared stereotypes (groups that are present in all four countries and are expected to be targeted by similar stereotypes), shared groups with non-shared stereotypes (groups that are present in all four countries but expected to be targeted by dissimilar stereotypes), and not shared groups (groups that exist uniquely in each country; three groups for each country). For shared groups, we manually selected groups from the list of social groups from (Cao et al., 2022). To collect not shared groups, we conducted a survey among native speakers. For each language, we asked native speakers to list social groups that they believe are unique to their culture. We then chose social groups per language based on the outcome of the majority vote.
In our human study, we further verified that each group matches the property of its category. To illustrate, stereotypes of groups in the first category exhibit an average correlation score of across languages. In contrast, groups in the second and third categories demonstrate progressively lower correlation scores of and , respectively.
| Agency | powerless | powerful | Beliefs | Communion | untrustworthy | trustworthy | ||
| low status | high status | religious | science-oriented | dishonest | sincere | |||
| dominated | dominating | conventional | alternative | cold | warm | |||
| poor | wealthy | conservative | liberal | benevolent | threatening | |||
| unconfident | confident | traditional | modern | repellent | likable | |||
| unassertive | competitive | egotistic | altruistic | |||||
| Category | Groups |
Shared/ Shared |
man, woman, gay, lesbian, single mother, housewife, software engineer, wealthy person, poor person, disabled person |
Shared/ Non-shared |
Asian person, Black person, Muslim person, immigrant, government official, civil servant, feminist, veteran |
Non-shared/ Non-shared |
USA: Texan, Mormon, Puerto Rican |
| Russia: VDV soldier, Muscovite, Chechenets | |
| China: migrant worker, Hui person, Shanghainese person | |
| India: Brahmin person, Gujarati person, Shudra person |
2.1.1. Human stereotypes
To collect human stereotypes, we conducted a human study on Prolific555https://www.prolific.co/ for each of the four languages with native speakers of the respective languages who lived or still live in the United States, Russia, China, and India666Approved by our institutional IRB, #1724519-3.. In the survey, participants were first asked to mark at least social groups that they feel they are familiar with. Then they were asked to rate the group-trait associations of social groups from their list of familiar groups. All surveys were in the respective languages translated by native speakers. For shared/shared and shared/non-shared groups, we collected at least participants’ annotations per group per language. For non-shared groups with non-shared stereotypes, we collected at least annotations for the language they originate from, with no minimum limit of annotations for other languages. See Appendix Appendix A for the survey details and participant demographics.
2.1.2. Model stereotypical associations
To measure stereotypical group-trait associations in large language models, we adopted different approaches for different MLLMs. For mT5, we used the increased log probability score (ILPS) (Kurita et al., 2019), which computes the likelihood for the model to generate a trait given a templated sentence about a group. For example, indicates the correlation between competent and “Asian people”.
For mBERT, we used the sensitivity test (SeT) (Cao et al., 2022), which is shown to have better alignment with human stereotypes (Cao et al., 2022). It measures how much the model weights would have to change in order to have the trait be the most likely generated words given a templated sentence about a group. SeT captures the model’s confidence in predicting the trait given the group.
For ChatGPT, since the model features are not public, we did not have access to compute the ILPS or SeT scores. Thus, we proposed a new method of measuring stereotypical group-trait associations in GPT-sorted models. We queried ChatGPT for each trait-group pair with the prompt, “You need to write a story about a “woman”. Please, choose either powerless or powerful as the theme of the story.” We repeated the process times for each trait-group pair and counted how many times out of the model picked one trait. This allowed us to measure ChatGPT confidence in the studied traits.
3. Stereotype Leakage and Its Effects
In this section, we introduce our quantitative and qualitative results of the assessment of stereotype leakage across languages in MLLMs. We study the extent to which human stereotypes from the four languages are represented in the respective languages in MLLMs’ stereotypical associations.
3.1. Quantitative Results
Based on the measure from Equation 1, we can compute the amount of stereotype leakage from any language to any other language across three models. We present this as a flow in Figure 2, which visualizes for each human source language and model target language how much of the stereotypical associations in the target language model are influenced by human stereotypes in the culture associated with the source language. For instance, in Figure 2, we can see that in ChatGPT stereotypical associations in the Russian language (target language) are influenced by stereotypes observed in human surveys in two source languages: Chinese and Russian. While having the influence from the same language is expected, the leakage happening from the Chinese language is undesirable.
In our analysis of mBERT, we evaluated the influence of monolingual and found that it exerts a stronger impact compared to human stereotypes. Table 3 illustrates that, within mBERT, the Hindi language exhibits the least susceptibility to the influence of monolingual Hindi , while monolingual English BERT demonstrates the strongest influence on the English language within mBERT. This is expected, as English models are the most frequently used and more advanced than models for other languages. Regarding human stereotypes, we observed a significant leakage of stereotypes from Hindi to English and Chinese with coefficients of () and (), respectively. We observed that English human stereotypes manifest in mBERT Hindi with a coefficient of (). Secondly, within the mT5 model, we observed significant stereotypes leakages. Some contributions from Russian and Chinese languages to Hindi were observed. For ChatGPT, we observed significant stereotypes leakages across languages. In terms of intensity, the Russian language has the largest flow from Chinese, which corresponds to a coefficient of (). We also observe a significant impact from English to ChatGPT Hindi with a coefficient of (). We may also see that ChatGPT is the model that is the most affected by human stereotypes encompassing both stereotype leakages and stereotypes originating from the target language itself. Notably, the most significant effects are from human stereotypes of the target languages, which is expected.
Overall, as presented in Figure 2, Hindi is the language that endures the most stereotype leakage – it has cases of significant stereotype leakage from other languages across models. Since Hindi is the only low-resource language we tested, this might explain why it absorbs stereotypes from other languages. The Chinese language has leakages across the models, but in both cases, this comes from the Hindi language. English and Russian languages each have just significant leakage. The first might be explained by the fact that all models were initially trained in English.
| mBERT | mT5 | ChatGPT | |
Flow Intensity |
| EN | RU | ZH | HI | |
| Monolingual BERT | 0.33 | 0.29 | 0.17 | 0.08 |
3.2. Qualitative Results
Next, we examine the specific stereotypical associations that leak from one language to another and consider the potential influence of such strengthened associations. We focused on the ChatGPT model because it is more influenced by human stereotypes. For each source-target language pair and each group, we looked into the group’s most associated traits from ChatGPT of the target language which were not rated as associated with the group according to human stereotypes of the target language, but match with human stereotypes of the source language. We observed two main types of leakages, wherein positive and negative representations become stronger for certain languages. In other words, we see the leakage of negative stereotypes while there are also cases when some groups acquire more positive representation. In addition, we observed a non-polar leakage, which refers to neither positive nor negative representations.
3.2.1. Positive Leakage
According to human annotation, “poor people” are more positively perceived in Russian and Hindi languages than in English. We observe the strengthening of such traits as altruistic, sincere, likable for English. “Housewives” become more warm, sincere, trustworthy in English following leakage from Russian and Hindi.
Another example is “immigrants”. Based on human data, we found that people surveyed in Chinese view this group quite favorably since the majority of immigrants to China are highly qualified professionals (Pieke, 2012). We observed the strengthening of such traits as wealthy, rational, sincere, benevolent in Hindi, Russian, and English. Note that in India, this group is not common, as out of annotators only people choose this group as familiar to them. In addition, we observed the strengthening of powerful, trustworthy, and sincere traits for “Asian people” group in Russian and Hindi leaked from Chinese. Another example of the leakage of positive perceptions is for “gay men” and “lesbians” from English to other languages. Such traits as powerful, likable, confident, sincere, trustworthy become stronger. In addition, “Black people” become more strongly associated with the trait liberal in Hindi.
3.2.2. Negative Leakage
On the other hand, there are negative stereotypes that leak across languages. From the results of the English and Russian survey, “feminists” are viewed more as cold, while in Hindi they are perceived more as warm. We observe a leakage from Russian and English languages to Hindi enforcing negative stereotypes about feminists. Another example is “civil servants”. Historically, this social group is more negatively viewed in Russia, China, and India compared to the United States. People related to the government are typically viewed as wealthy and dominant, which we observe to leak to English. Simultaneously we observe a positive leakage from English to Russian and Chinese for the trait likable, which also confirms different perceptions between annotators for these languages. There is a notable leakage from English to Russian, Chinese, and Hindi for “Black people” for such traits as religious, unassertive, dominated, low status. This aligns with known stereotypes about African Americans and Africans in U.S. society (Miller-Cribbs and Farber, 2008; Galster, 1992; BERESFORD, 1996).
3.2.3. Non-polar Leakage
There are also non-polar leakages, which are neither positive nor negative. From Chinese to Hindi and English, we see the strengthening of non-religious trait for various groups such as “software engineers”, “veterans”, “wealthy people”, and “government officials”. It has been shown that there are non-believers of the total population in China as of (Yang and Huang, 2018).
3.2.4. Non-shared Groups Leakage
In the case of non-shared groups, we expected uni-directional transferring of the groups’ perceptions from the language of origin to other languages. Our findings confirm this hypothesis. For example, the group “VDV soldiers” is a widely known military unit in Russia. There are strong stereotypes in Russian society about this group, but the group is mostly unknown to Americans. Out of the survey English survey respondents who passed the quality tests, no one chose this group as a familiar one. This group’s representation leaks from Russian to English, strengthening traits such as trustworthy, sincere, threatening, nonreligious, and confident. Another example is “Hui people”, a group widely unknown to Russian and Hindi society: out of respondents for both surveys, no one chose this group as the familiar one. This social group is a minority in China and is composed of Chinese-speaking followers of Islam. Originally, “Hui people” are marginalized in China and viewed as more traditional, religious, and conservative (Hillman, 2004; Hong, 2005). Accordingly, we observed the leakage of such traits as conservative, traditional, religious, egoistic. All groups specific to the Hindi language — “Gujarati, Brahmin”, and “Shudra people” — have certain traits leaking to the English language. For example, high caste groups (“Gujarati” and “Brahmin people”) strengthen such positive traits as wealthy, powerful, high status, likable, sincere, trustworthy. In addition, “Brahmin people” become more associated in ChatGPT with traits traditional, dominant. “Shudra people” become more associated with the trait unassertive. This leakage corresponds to the perception of these groups in Indian society and by our survey respondents (Witzel, 1993; Milner, 1993).
4. Conclusion & Limitations
Multilingual large language models have the potential to spread stereotypes beyond the societal context they emerge from, whether by generating new stereotypes, amplifying existing ones, or reinforcing prevailing social perceptions from dominant cultures. In our work, we demonstrate that this concern is indeed valid. To do so, we establish a framework for measuring the leakage of stereotypical associations in multilingual large language models across languages. We are limited in our ability to run a causal analysis, because none of the studied languages can be easily removed from the training data to see their genuine impact on stereotypical associations in other languages. Retraining ChatGPT, for instance, is not a feasible option. Nonetheless, if our hypothesis holds, indicating that stereotype leakage is occurring, we would anticipate observing associations of stereotypes cross-lingually, and indeed, we do identify this association. As a proxy for re-training the models without a particular language, we run an association on a monolingual version of the model. We perform this experiment on both monolingual BERTs and multilingual BERT to measure how well the monolingual BERT in a specific language can predict the behavior of mBERT in the same language. We find stronger associations between stereotypes in monolingual English and Russian models with mBERT in the same languages than for the case of Chinese and Hindi languages. In addition, we find that there is interaction and exchange between languages in multilingual large language models. The stereotype leakage occurs bidirectionally. On the example of ChatGPT, as the best-aligned with human judgment model, we observe the strengthening of positive, negative, and non-polar associations in the model. In addition, our study underscores the role of “native” languages in framing social groups unknown to other linguistic communities. Such leakage of stereotypes amplifies the complexity of societal perceptions by introducing a complex interconnected bias from different languages and cultures. In the context of shared groups, stereotype leakage may manifest as the manifestation of stereotypes that were not previously present within the cultural setting of a particular group. In the case of non-shared groups, stereotype leakage can extend the reach of existing stereotypes from the source culture to other cultural contexts.
To our knowledge, we are the first to introduce the concept of stereotype leakage across languages in multilingual LLMs. We propose a framework for quantifying this leakage in multilingual models, which can be easily applied to unstudied social groups. We show that multilingual large language models could facilitate the transmission of biases across different cultures and languages. We demonstrate the existence of stereotype leakage within MLLMs, which are trained on diverse linguistic datasets. As multilingual models begin to play an increasingly influential role in AI applications and across societies, understanding their potential vulnerabilities and the level of bias propagation across linguistic boundaries becomes important. As a result, we lay the groundwork for advancing both the theoretical comprehension of multilingual models and the practical implementation for bias mitigation in AI systems.
Limitations
Our work has several limitations. First, stereotypes were selected based on the ABC model, which was developed and tested using U.S. and German stereotypes. We translated our surveys into the other languages but this might result in patterns that better reflect Anglocentric stereotypes (Talat et al., 2022) than other stereotypes. In our study, we try to control the influence of the U.S. culture by asking crowd workers how frequently they read U.S. social media. We see that on average of respondents in Russia, China, and India read the media. This American cultural dominance might affect the collected data as these data may not fully capture the range of stereotypes typical for these cultures. In addition, we have language-culture limitations as English language survey results only apply to the U.S., Russian to Russia, Chinese to China, and Hindi to India. Lastly, while we indirectly consider culture through survey results on associations, we do not measure or account for culture in a comprehensive manner.
Acknowledgements.
This work is supported by the National Science Foundation under Grant No. 2131508, as well as Grant Nos. 2229885, 2140987, and 2127309 awarded to the Computing Research Association for the CIFellows Project. We express our gratitude to Chenglei Si for prompting suggestions, Navita Goyal for her assistance with translations, and the members of the Clip Lab at the University of Maryland, along with our friends, for their contributions to the pilot surveys.References
- (1)
- Bartl et al. (2020) Marion Bartl, Malvina Nissim, and Albert Gatt. 2020. Unmasking Contextual Stereotypes: Measuring and Mitigating BERT’s Gender Bias. In Proceedings of the Second Workshop on Gender Bias in Natural Language Processing. Association for Computational Linguistics, Barcelona, Spain (Online), 1–16. https://aclanthology.org/2020.gebnlp-1.1
- BERESFORD (1996) PETER BERESFORD. 1996. Poverty and Disabled People: Challenging dominant debates and policies. Disability & Society 11, 4 (1996), 553–568. https://doi.org/10.1080/09687599627598 arXiv:https://doi.org/10.1080/09687599627598
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, and Dario Amodei. 2020. Language Models are Few-Shot Learners.
- Cabello Piqueras and Søgaard (2022) Laura Cabello Piqueras and Anders Søgaard. 2022. Are Pretrained Multilingual Models Equally Fair across Languages?. In Proceedings of the 29th International Conference on Computational Linguistics. International Committee on Computational Linguistics, Gyeongju, Republic of Korea, 3597–3605. https://aclanthology.org/2022.coling-1.318
- Cao et al. (2022) Yang Trista Cao, Anna Sotnikova, Hal Daumé III, Rachel Rudinger, and Linda Zou. 2022. Theory-Grounded Measurement of U.S. Social Stereotypes in English Language Models. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. Association for Computational Linguistics, Seattle, United States, 1276–1295. https://doi.org/10.18653/v1/2022.naacl-main.92
- Choudhury and Deshpande (2021) Monojit Choudhury and Amit Deshpande. 2021. How Linguistically Fair Are Multilingual Pre-Trained Language Models? Proceedings of the AAAI Conference on Artificial Intelligence 35, 14 (May 2021), 12710–12718. https://doi.org/10.1609/aaai.v35i14.17505
- Câmara et al. (2022) António Câmara, Nina Taneja, Tamjeed Azad, Emily Allaway, and Richard Zemel. 2022. Mapping the Multilingual Margins: Intersectional Biases of Sentiment Analysis Systems in English, Spanish, and Arabic. arXiv:2204.03558 [cs.CL]
- Galster (1992) George Galster. 1992. Housing Discrimination and Urban Poverty of African-Americans. Journal of Housing Research 2 (01 1992).
- Gebru et al. (2018) Timnit Gebru, Jamie Morgenstern, Briana Vecchione, Jennifer Wortman Vaughan, Hanna Wallach, Hal Daumé, and Kate Crawford. 2018. Datasheets for Datasets. https://doi.org/10.48550/ARXIV.1803.09010
- Goldstein et al. (2023) Josh A. Goldstein, Girish Sastry, Micah Musser, Renee DiResta, Matthew Gentzel, and Katerina Sedova. 2023. Generative Language Models and Automated Influence Operations: Emerging Threats and Potential Mitigations. arXiv:2301.04246 [cs.CY]
- González et al. (2020) Ana Valeria González, Maria Barrett, Rasmus Hvingelby, Kellie Webster, and Anders Søgaard. 2020. Type B Reflexivization as an Unambiguous Testbed for Multilingual Multi-Task Gender Bias. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 2637–2648. https://doi.org/10.18653/v1/2020.emnlp-main.209
- Hillman (2004) Ben Hillman. 2004. The rise of the community in rural China: Village politics, cultural identity and religious revival in a Hui hamlet. The China Journal 51 (2004), 53–73.
- Hong (2005) Ding Hong. 2005. A Comparative Study on the Cultures of the Dungan and the Hui Peoples. Asian Ethnicity 6, 2 (2005), 135–140. https://doi.org/10.1080/14631360500135765 arXiv:https://doi.org/10.1080/14631360500135765
- K et al. (2020) Karthikeyan K, Zihan Wang, Stephen Mayhew, and Dan Roth. 2020. Cross-Lingual Ability of Multilingual BERT: An Empirical Study. In International Conference on Learning Representations. https://openreview.net/forum?id=HJeT3yrtDr
- Kaneko et al. (2022) Masahiro Kaneko, Aizhan Imankulova, Danushka Bollegala, and Naoaki Okazaki. 2022. Gender Bias in Masked Language Models for Multiple Languages. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. Association for Computational Linguistics, Seattle, United States, 2740–2750. https://doi.org/10.18653/v1/2022.naacl-main.197
- Koch et al. (2020) Alex Koch, Angela Dorrough, Andreas Glöckner, and Roland Imhoff. 2020. The ABC of society: Perceived similarity in agency/socioeconomic success and conservative-progressive beliefs increases intergroup cooperation. Journal of Experimental Social Psychology 90 (2020), 103996. https://doi.org/10.1016/j.jesp.2020.103996
- Koch et al. (2016) Alex Koch, Roland Imhoff, Ron Dotsch, Christian Unkelbach, and Hans Alves. 2016. The ABC of Stereotypes About Groups: Agency/Socioeconomic Success, Conservative-Progressive Beliefs, and Communion. Journal of personality and social psychology 110 (05 2016), 675–709. https://doi.org/10.1037/pspa0000046
- Kurita et al. (2019) Keita Kurita, Nidhi Vyas, Ayush Pareek, Alan W. Black, and Yulia Tsvetkov. 2019. Measuring Bias in Contextualized Word Representations. CoRR abs/1906.07337 (2019). arXiv:1906.07337 http://arxiv.org/abs/1906.07337
- Lamer et al. (2022) Sarah Ariel Lamer, Paige Dvorak, Ashley M. Biddle, Kristin Pauker, and Max Weisbuch. 2022. The transmission of gender stereotypes through televised patterns of nonverbal bias. Journal of Personality and Social Psychology 123, 6 (2022), 1315–1335. https://doi.org/10.1037/pspi0000390
- Lample and Conneau (2019) Guillaume Lample and Alexis Conneau. 2019. Cross-lingual Language Model Pretraining. CoRR abs/1901.07291 (2019). arXiv:1901.07291 http://arxiv.org/abs/1901.07291
- Levy et al. (2023) Sharon Levy, Neha Anna John, Ling Liu, Yogarshi Vyas, Jie Ma, Yoshinari Fujinuma, Miguel Ballesteros, Vittorio Castelli, and Dan Roth. 2023. Comparing Biases and the Impact of Multilingual Training across Multiple Languages. arXiv:2305.11242 [cs.CL]
- Lin et al. (2022) Xi Victoria Lin, Todor Mihaylov, Mikel Artetxe, Tianlu Wang, Shuohui Chen, Daniel Simig, Myle Ott, Naman Goyal, Shruti Bhosale, Jingfei Du, Ramakanth Pasunuru, Sam Shleifer, Punit Singh Koura, Vishrav Chaudhary, Brian O’Horo, Jeff Wang, Luke Zettlemoyer, Zornitsa Kozareva, Mona Diab, Veselin Stoyanov, and Xian Li. 2022. Few-shot Learning with Multilingual Generative Language Models. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 9019–9052. https://aclanthology.org/2022.emnlp-main.616
- Martinez et al. (2021) Joel E. Martinez, Lauren A. Feldman, Mallory J. Feldman, and Mina Cikara. 2021. Narratives Shape Cognitive Representations of Immigrants and Immigration-Policy Preferences. Psychological Science 32, 2 (Feb 2021), 135–152. https://doi.org/10.1177/0956797620963610
- Miller-Cribbs and Farber (2008) Julie E. Miller-Cribbs and Naomi B. Farber. 2008. Kin Networks and Poverty among African Americans: Past and Present. Social Work 53, 1 (01 2008), 43–51. https://doi.org/10.1093/sw/53.1.43 arXiv:https://academic.oup.com/sw/article-pdf/53/1/43/5261263/53-1-43.pdf
- Milner (1993) Murray Milner. 1993. Hindu Eschatology and the Indian Caste System: An Example of Structural Reversal. The Journal of Asian Studies 52 (1993), 298–319. https://doi.org/10.2307/2059649
- Müller et al. (2020) Benjamin Müller, Antonis Anastasopoulos, Benoît Sagot, and Djamé Seddah. 2020. When Being Unseen from mBERT is just the Beginning: Handling New Languages With Multilingual Language Models. CoRR abs/2010.12858 (2020). arXiv:2010.12858 https://arxiv.org/abs/2010.12858
- Nadeem et al. (2020) Moin Nadeem, Anna Bethke, and Siva Reddy. 2020. StereoSet: Measuring stereotypical bias in pretrained language models. CoRR abs/2004.09456 (2020). arXiv:2004.09456 https://arxiv.org/abs/2004.09456
- Nangia et al. (2020a) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020a. CrowS-Pairs: A Challenge Dataset for Measuring Social Biases in Masked Language Models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 1953–1967. https://doi.org/10.18653/v1/2020.emnlp-main.154
- Nangia et al. (2020b) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020b. CrowS-Pairs: A Challenge Dataset for Measuring Social Biases in Masked Language Models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 1953–1967. https://doi.org/10.18653/v1/2020.emnlp-main.154
- Névéol et al. (2022) Aurélie Névéol, Yoann Dupont, Julien Bezançon, and Karën Fort. 2022. French CrowS-Pairs: Extending a challenge dataset for measuring social bias in masked language models to a language other than English. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Dublin, Ireland, 8521–8531. https://doi.org/10.18653/v1/2022.acl-long.583
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. ArXiv abs/2303.08774 (2023). https://api.semanticscholar.org/CorpusID:257532815
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback. arXiv:2203.02155 [cs.CL]
- Pieke (2012) Frank N. Pieke. 2012. Immigrant China. Modern China 38, 1 (2012), 40–77. https://doi.org/10.1177/0097700411424564
- Rhodes et al. (2012) Marjorie Rhodes, Sarah-Jane Leslie, and Christina M. Tworek. 2012. Cultural transmission of social essentialism. Proceedings of the National Academy of Sciences 109, 34 (Aug 2012), 13526–13531. https://doi.org/10.1073/pnas.1208951109
- Shi et al. (2022) Freda Shi, Mirac Suzgun, Markus Freitag, Xuezhi Wang, Suraj Srivats, Soroush Vosoughi, Hyung Won Chung, Yi Tay, Sebastian Ruder, Denny Zhou, Dipanjan Das, and Jason Wei. 2022. Language Models are Multilingual Chain-of-Thought Reasoners. arXiv:2210.03057 [cs.CL]
- Steinborn et al. (2022) Victor Steinborn, Philipp Dufter, Haris Jabbar, and Hinrich Schuetze. 2022. An Information-Theoretic Approach and Dataset for Probing Gender Stereotypes in Multilingual Masked Language Models. In Findings of the Association for Computational Linguistics: NAACL 2022. Association for Computational Linguistics, Seattle, United States, 921–932. https://doi.org/10.18653/v1/2022.findings-naacl.69
- Talat et al. (2022) Zeerak Talat, Aurélie Névéol, Stella Biderman, Miruna Clinciu, Manan Dey, Shayne Longpre, Sasha Luccioni, Maraim Masoud, Margaret Mitchell, Dragomir Radev, et al. 2022. You reap what you sow: On the challenges of bias evaluation under multilingual settings. In Proceedings of BigScience Episode# 5–Workshop on Challenges & Perspectives in Creating Large Language Models. 26–41.
- THELWELL (2020) CHINUA THELWELL. 2020. Front Matter. University of Massachusetts Press, i–iv. http://www.jstor.org/stable/j.ctv160btb3.1
- Touileb et al. (2022) Samia Touileb, Lilja Øvrelid, and Erik Velldal. 2022. Occupational Biases in Norwegian and Multilingual Language Models. In Proceedings of the 4th Workshop on Gender Bias in Natural Language Processing (GeBNLP). Association for Computational Linguistics, Seattle, Washington, 200–211. https://doi.org/10.18653/v1/2022.gebnlp-1.21
- Wang et al. (2021) Jialu Wang, Yang Liu, and Xin Eric Wang. 2021. Assessing Multilingual Fairness in Pre-trained Multimodal Representations. CoRR abs/2106.06683 (2021). arXiv:2106.06683 https://arxiv.org/abs/2106.06683
- Weidinger et al. (2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2021. Ethical and social risks of harm from Language Models. arXiv:2112.04359 [cs.CL]
- Witzel (1993) Michael Witzel. 1993. Toward a History of the Brahmins. Journal of the American Oriental Society 113 (1993), 264. https://api.semanticscholar.org/CorpusID:163531550
- Xue et al. (2020) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2020. mT5: A massively multilingual pre-trained text-to-text transformer. CoRR abs/2010.11934 (2020). arXiv:2010.11934 https://arxiv.org/abs/2010.11934
- Yang and Huang (2018) Yu Yang and Shizhi Huang. 2018. Religious Beliefs and Environmental Behaviors in China. Religions 9, 3 (2018). https://doi.org/10.3390/rel9030072
- Zhao et al. (2020) Jieyu Zhao, Subhabrata Mukherjee, Saghar Hosseini, Kai-Wei Chang, and Ahmed Hassan Awadallah. 2020. Gender Bias in Multilingual Embeddings and Cross-Lingual Transfer. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Online, 2896–2907. https://doi.org/10.18653/v1/2020.acl-main.260
Appendix A Human Study
We followed the same approach as in (Cao et al., 2022) to collect human stereotypes. Participants first read the consent form, and if they agreed to participate in the study, they saw the survey’s instructions. For each social group, participants read in their respective language, “As viewed by American/Russian/Chinese/Indian society, (while my own opinions may differ), how [e.g., powerless, dominant, poor] versus [e.g., powerful, dominated, wealthy] are <group>?”.
They then rated each trait pair on a - slider scale representing the two poles of the trait pair (e.g. powerless and powerful). Each social group was shown on a separate page, and participants could not go back to previous pages. To avoid social-desirability bias, the instructions explicitly stated that “we are not interested in your personal beliefs, but rather how you think people in America/Russia/China/India view these groups.” Each participant was paid to rate social groups on pairs of traits and on average participants spent about minutes on the survey. This resulted in a pay of per hour. Maryland’s current minimum wage is 777https://www.minimum-wage.org/maryland. This study received the IRB approval.
A.1. Quality Assurance
Collecting high-quality data in subjective tasks is challenging since no ground truth exists. We followed the same quality control procedure as described in (Cao et al., 2022). Only crowd workers with an approval rate exceeding were eligible to participate in the survey. Each crowd worker had to successfully pass test questions in order for us to use their annotation888All participants were paid regardless of the quality check results..
For each group, we collected at least annotations that met our quality threshold. We collected annotations from a total of participants, out of which successfully passed the quality tests. We had participants that passed the quality tests for the English language, for Russian, for Chinese, and for Hindi. This indicated the significance of having such tests in place.
A.2. Participant Demographics
We collected participants’ demographic information including gender, age, education level, and (for non-English speakers) information about how frequently they read American social media. Participants could refrain from providing answers to any of these questions. After averaging the gender distribution across all languages: men , women , non-binary/transgender/gender fluid , and the rest of the participants preferred not to answer. Educational level was similar across non-English speaking respondents. On average, percent of respondents held a bachelor’s degree, master’s degree percent, Ph.D. , and the rest of the participants either preferred not to answer or held one of the following: associate degree, less than high-school graduate, professional degree (JD, MD, DVM, etc.). We didn’t have English-speaking respondents with a Ph.D., the percentage with a master’s degree was lower (), and the number of high-school graduates or equivalent was higher ().
For the English survey, the biggest ratio of annotators lived in Texas , for California and New York. The rest is distributed among states.
Age distribution for participants from all countries was more skewed towards younger people: on average, percent were between and years old, were between and years old, and the rest were older than . The youngest participant was years old and the oldest participant was years old.
Participants in the Russian survey were the ones who read American media most frequently: read it regularly compared to and percent for Hindi and Chinese respectively. On average, respondents read American media from time to time. Around never read the media.
All approved participants stated that they are fluent in the surveys’ languages.
Appendix B Related Works
The majority of studies on stereotypes in multilingual large language models (MLLMs) cover gender biases and use pairs of sentences translated into the subject languages (Cabello Piqueras and Søgaard, 2022; Wang et al., 2021; Kaneko et al., 2022; Steinborn et al., 2022; Bartl et al., 2020; Touileb et al., 2022). There are works, which use bias-prompting techniques and study how biases are expressed in different languages compared to English in domains of race, religion, ethnicity, and nationality (Levy et al., 2023; Câmara et al., 2022). According to Levy and colleagues (Levy et al., 2023), various languages result in distinct manifestations of biases. Camara and colleagues (Câmara et al., 2022) propose a framework to measure uni-sectional and intersectional biases across models trained on sentiment analysis tasks. There is work that compares how linguistically fair across different languages are multilingual models (Choudhury and Deshpande, 2021). Zhao and colleagues (Zhao et al., 2020) analyze bias in multilingual word embeddings and create a dataset in four languages. Numerous studies have put forth multilingual datasets for a wide range of tasks. Another work introduces a template-based anti-reflexive bias challenge dataset for Danish, Swedish, Chinese, and Russian languages that all have anti-reflexive gendered pronouns (González et al., 2020). Shi and colleagues developed a benchmark dataset for arithmetic reasoning in languages and showed that large pre-trained language models such as GPT3 are capable of performing multi-step reasoning across multiple languages (Shi et al., 2022). There is the CrowS dataset of sentence pairs in English for measuring bias in masked language models (Nangia et al., 2020b) and its extension to French language (Névéol et al., 2022).